Use the graphing method to determine the answers of the following questions:

1- What is the order of the reaction?

2- Calculate the reaction rate with the appropriate units

3- Calculate the half-life t1/2 The data table is given below (carried at 430oC)

C4H8 (g) → 2C2H4 (g)

Time (s) Pressure C4H8 (mmHg)
0 400
2000 316
4000 248
6000 196
8000 155
10000 122
Answer #1

near FRof B 6.5 Badard Emor 59312 6.0 3 5.5 5.0 4.5 2000 0 2000 4000 6000 8000 10000 Time (s)

First Plot ln(Pressure) vs Time to get the plot using excel or origin software.

By inspection, the plot of ln(P) vs time is linear, and the plot of 1/Pa vs time is not.

When ln PC4H8 is plotted as a function of time, a straight line results.

- The process is first-order.

- k is the negative slope: 1.19 x 10-4

2.

Rate = -\DeltaPC4H8 / \Deltat = k PC4H8 = ()

3.

t1/2 for first order reaction is expressed as

t1/2 = 0.693 / k

= 0.693 / (1.19 x 10-4)

= 5.82 x 10-3 s